[Hallc_running] [New Logentry] RC update - July 24, 2025

arunin at jlab.org arunin at jlab.org
Thu Jul 24 10:00:01 EDT 2025


Logentry Text:
--
Data taking proceeded smoothly during both the swing and owl shifts. We began with the z = 0.5 π-SIDIS setting for the second LD2 cycle and continued operations with the solid targets. During the beam studies downtime, the N₂ flow into the beam dump was repaired, enabling us to reach beam currents of approximately 30 μA. Before resuming production after the beam studies, we requested a harp scan (https://logbooks.jlab.org/entry/4425188).

Yesterday, during the beam study period, Bill Hendry worked on addressing the cable reflection issue. While the peak at 120 ns after the main peak appears reduced, some reflections are still present (https://logbooks.jlab.org/entry/4425257).

Additionally, Ellen successfully obtained an HMS DP NMR lock using an oscilloscope and measured the magnetic field to be 0.418877 T on probe 3 (https://logbooks.jlab.org/entry/4425128).

We will hold our regular RC meeting today at 2:30 PM. Zoom link is provided below.


https://jlab-org.zoomgov.com/j/1618227115?pwd=8JkhFg6YNesVc3HNgWjwUSIKIu...

Meeting ID: 161 822 7115
Passcode: 253463


---

This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4425770
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20250724/a6e9f628/attachment.htm>


More information about the Hallc_running mailing list